OpenStudy (anonymous):

find complex zeros x^3-1

OpenStudy (anonymous):

graph it and see when the line encounter the x axis.

OpenStudy (solomonzelman):

\[x^3-1=0\]\[x^3-1\color{red} {^3}=0\]\[\color{red} {a^3-b^3}=\color{red} {(a-b)(a^2+ab+b^2)}\]\[\color{red} {x^3-1^3}=\color{red} {(x-1)(x^2+x+1)}\]\[\color{rblack} {(x-1)(x^2+x+1)}=0\]

OpenStudy (solomonzelman):

Solve for x.

OpenStudy (solomonzelman):

\[x-1=0~~~~~x=1~~~~OR,\]

OpenStudy (solomonzelman):

\[x^2+x+1=0\]\[\color{black}{\frac{1±i \sqrt{3}}{2}}\]\[\color{blue}{1}~~~,~~~\color{blue}{\frac{1+i \sqrt{3}}{2}}~~~and~~~\color{blue} {\frac{1-i \sqrt{3}}{2}}~~~\]are the zeros

OpenStudy (solomonzelman):

since the last 2 zeros are imaginary the graph would look like just x=1.

OpenStudy (anonymous):

solve for x